Ṅumbers, Booleans, strings; sets, functions, relations; digraphs, graphs, trees; state machinesṪheir properties. How to do them, how to read them, how to recognize correct/incorrect onesṘeasoning about propositions, about predicates (with quantifiers)Ċommon proof strategies: Direct proof, proof by contradiction, by casesİnduction (very important in this course and elsewhere in CS) ˙We’ll be working on this topic throughout the term. Basic concepts and methods underlying formal specification and verification of software systemsḊeveloping models for programs and their required behaviorṖroving that a program model satisfies its correctness requirements (6.170).